OpenStudy (anonymous):

what is radical sign 8 plus radical sign 72?

OpenStudy (anonymous):

2 radical sign 2 + 2radical sign 6

OpenStudy (unklerhaukus):

\[\sqrt8+\sqrt{72}\] \[8=2\times2\times2=2^3\]\[\sqrt{8}=\sqrt{2^3}=\sqrt{2^2\times 2}=2\sqrt2\] \[72=2\times6\times6=2\times6^2\] \[\dots\]\[\dots\]

OpenStudy (anonymous):

so whats the awnser

OpenStudy (unklerhaukus):

\[\sqrt{72}=\sqrt{2\times6^2}=6\sqrt2\]

OpenStudy (anonymous):

ok thanks!

OpenStudy (unklerhaukus):

can you add the two together by your self?

OpenStudy (anonymous):

i dont think so...

OpenStudy (unklerhaukus):

\[2\sqrt2+6\sqrt{2}=(2+6)\sqrt2=\cdots\]

OpenStudy (anonymous):

okay so \[\sqrt{72} + \sqrt{8}\] = ??

OpenStudy (unklerhaukus):

\[\sqrt{72}+\sqrt{8}=2\sqrt2+6\sqrt2=(6+2)\sqrt2=\dots\]

OpenStudy (anonymous):

okay im in advanced 7th grade do you think my teacher will be like curious as tyo why i do the problem so well??

OpenStudy (unklerhaukus):

did you get the final answer?

OpenStudy (anonymous):

well is the awnser what you said it's not like a one number awnser it's what you typed?

OpenStudy (unklerhaukus):

ok the final answer is \[\cdots=8\sqrt2\] I can explain any step you do not understand